# Do Schwartz Higher-Order Values Help Sentence-Level Human Value Detection? When Hard Gating Hurts

### v1

@@ 3.1 Problem formulation and label spaces
We study sentence-level human value detection under Schwartz’s refined theory (Schwartz, 2012). Each sentence may express none, one, or multiple values. Let be a sentence and the binary vector over the 19 basic values.111The benchmark provides attained and constrained annotations per value; we collapse them into a single expressed value signal (Section 3.2).
HO categories
To test whether coarser abstractions help, we deterministically derive eight HO binary labels from the 19 values following Schwartz (2012): Openness to Change, Conservation, Personal Focus, Social Focus, Self-Enhancement, Self-Transcendence, Growth, and Self-Protection. Let be the set of HO categories and the basic values grouped under . We define:
@@
Presence label
@@
Bipolar evaluation slices
In addition to performance over all eight HO labels, we analyze the four canonical bipolar pairs to expose asymmetries in learnability and error propagation: (i) Openness to Change vs. Conservation, (ii) Self-Enhancement vs. Self-Transcendence, (iii) Personal Focus vs. Social Focus, and (iv) Growth vs. Self-Protection (Schwartz, 2012). For each pair, we report Macro- averaged over the two poles.
@@ 3.2 Dataset and preprocessing
We use the official train/validation/test split (The ValuesML Team, 2024; Touché, 2024) at sentence level. To keep modeling choices unified and compute controlled, we use the benchmark’s English version (machine-translated sentences provided by the benchmark). This yields sentences: train, validation, and test. The split is at the text level; all sentences inherit their source text’s split.
@@ 3.3.1 Direct multi-label prediction (supervised encoder)
The Direct approach follows Yeste and Rosso (2026). Given the pooled sentence representation , we apply a linear layer to produce logits and probabilities . Training minimizes standard multi-label binary cross-entropy: we minimize standard multi-label binary cross-entropy over labels. We do not use class weights, matching Yeste and Rosso (2026).
@@ 3.3.3 PresenceCategoryValues cascade
We test a three-stage cascade where Presence acts as a first gate. The Presence formulation follows Yeste and Rosso (2026); the novelty here is the PresenceHOValues cascade:
@@
This cascade can improve precision by suppressing spurious positives on non-value sentences, but it can compound errors across stages. We quantify overall and per-pair effects in Section 3.8.
@@ 3.3.4 Instruction-tuned LLM baselines (definition prompting)
We benchmark instruction-tuned open LLMs that fit on a single 8 GB GPU (Llama 3.1 8B, Ministral 8B 2410, Qwen 2.5 7B, Gemma 2 9B). We use the definition-style prompt (best in Yeste and Rosso (2026)), presenting one-line definitions for the 19 values from Schwartz (2012). The model returns only a JSON array of applicable value names.
We evaluate zero-shot and few-shot prompting with in-context examples. Few-shot templates prepend exemplars in the same schema and include at least one null exemplar (empty array) when . We use greedy decoding with max_new_tokens=200.
Post-processing and label mapping
@@ 3.3.5 QLoRA fine-tuning (parameter-efficient LLM adaptation)
We additionally evaluate supervised QLoRA (Dettmers et al., 2023), following Yeste and Rosso (2026). Based on validation screening, we fine-tune Gemma 2 9B as the backbone. We train only low-rank adapters (base model frozen) with learning rate and save only adapter weights.
@@ 3.4 Compute-frugal auxiliary signals
Lexicon-derived features. We build sentence vectors from psycholinguistic, affective, moral, and value resources: LIWC-22 (Boyd et al., 2022), eMFD (Hopp et al., 2021), the Schwartz value lexicon in ValuesML (Kiesel et al., 2023), and affective lexica including NRC VAD (Warriner et al., 2013; Mohammad, 2018), NRC EmoLex (Mohammad and Turney, 2013), NRC Emotion Intensity (Mohammad and Kiritchenko, 2018), and WorryWords (Mohammad, 2024). We aggregate token signals into sentence statistics (counts, relative frequencies, averaged intensities) and standardize using training-set statistics.
@@ 3.5 Training protocol and compute parity
Unless otherwise stated, we follow the supervised protocol of Yeste and Rosso (2026). All transformer models fine-tune microsoft/deberta-base (He et al., 2021) with a linear multi-label head. Inputs are tokenized and truncated/padded to length 512. We optimize with AdamW (Loshchilov and Hutter, 2019) and a linear schedule with warmup, using batch size 4, gradient accumulation 4 (effective batch size 16), learning rate , weight decay 0.15, and up to 10 epochs with early stopping on validation Macro- (patience 4). Dropout is 0.1.
@@
To keep the study compute-frugal while covering many strategies (direct prediction, hard gates/cascades, auxiliary signals, prompting, QLoRA, ensembles), we fix a single random seed (as in Yeste and Rosso (2026)) for supervised runs. This prioritizes breadth under a fixed budget and keeps differences tied to modeling choices rather than to random initialization. Instead of multiple seeds, we emphasize paired evaluation: nonparametric sentence-level bootstrap uncertainty for Macro- and per-label paired tests (McNemar with FDR correction). We interpret differences below roughly 1–2 Macro- points conservatively and focus on effects supported by paired tests.
@@ 3.6 Threshold calibration
For supervised encoders (and QLoRA models that output probabilities), we use (i) a fixed global threshold or (ii) label-wise thresholds from a constrained grid search over . For each label , we maximize validation recall subject to precision (ties broken by higher recall). For hierarchical models, thresholds are tuned in stage-aware order (Presence first, then HO, then values). Prompted LLMs output discrete label sets and do not require threshold calibration.
@@ 3.8 Evaluation metrics and statistical testing
Primary metric
We report macro-averaged over the label set. We compute per label across all sentences, then average across labels. For bipolar slices (Section 3.1), we compute Macro- for each pole and average the two poles.
End-task evaluation
@@
Uncertainty and significance
To assess robustness, we use nonparametric bootstrap resampling over sentences (Efron, 1992). We draw samples with replacement, recompute Macro-, and estimate (i) a one-sided 95% lower bound for and (ii) a one-sided empirical -value.
Per-label paired tests
@@ 3.9 Reproducibility
We log all configurations (preprocessing, model variants, thresholds, ensemble membership) and preserve prediction files for all runs. We also release the value-to-HO mapping and scripts for preprocessing and evaluation to enable exact replication on the benchmark splits (Touché, 2024). For LLM experiments, we release the prompts/templates and any adapter weights (where redistribution is permitted), along with deterministic decoding settings and post-processing scripts.
@@ 4.1 Higher-order categories are learnable, but not equally so
HO categories are learnable with compact supervised encoders, but difficulty varies by pair. The easiest pair is Growth vs. Self-Protection (test Macro-, Table S3). Self-Transcendence vs. Self-Enhancement is moderate (best Macro-, Table S9), while Openness vs. Conservation is hardest (best Macro-, Table S7). These patterns track label prevalence: Openness is rare (about 8% vs. about 20% for Conservation, Table 2), which likely drives both lower Macro- and strong pole asymmetry (Openness Conservation in Table 1).
Table 1 summarizes these differences: Growth vs. Self-Protection is most learnable (Macro-), Self-Transcendence vs. Self-Enhancement is moderate (best Macro-), and Openness vs. Conservation remains hardest (best Macro-), with persistent asymmetry (Conservation Openness). The pattern suggests that “constraint/tradition” cues are captured more reliably than “novelty/autonomy” cues at sentence level. We report single-run results and interpret differences under 1–2 Macro- points cautiously.
@@ 4.4 Presence gating boosts in-gate validation scores but does not robustly improve end-task performance
The most plausible explanation is error compounding: Presence false negatives suppress downstream positives (recall loss), while Presence false positives admit sentences that still look negative to the HO classifier (precision loss). The net effect is that hard gating trades recall for precision in a way that is hard to tune globally. This sensitivity is visible empirically: changing the gate threshold (e.g., 0.50 vs. 0.10) often leaves test Macro- unchanged or slightly worse (Tables S11 and S13–S17). This motivates softer alternatives rather than a binary filter.
@@ 4.5 HOvalues hard gating does not translate into reliable gains on fine-grained detection
Under the Growth gate, the best test Macro- for Growth values is about ; under Self-Protection it is about . Section S7 compares the best HO-gated systems against tuned Direct baselines within each HO slice using paired bootstrap tests. The pattern is consistent: hard HO gating does not yield reliable improvements and is sometimes significantly worse, consistent with recall suppression when parent categories are missed. In a sentence-level setting—where signals are short, implicit, and often multi-valued—this loss is hard to recover.
@@ 4.6 Where small instruction-tuned LLMs fit under the same budget
Section S4 benchmarks instruction-tuned 10B LLMs (prompted and QLoRA-adapted) on the same HO-restricted slices. Table 4 reports the best test Macro- for Gemma-2-9B-it under (i) prompting only, (ii) prompting plus a lightweight SBERT gate, and (iii) QLoRA adaptation, and indicates which upgrades are supported by bootstrap tests (Section S7). Overall, even the best LLM variants remain well below the supervised DeBERTa-based champions on test (e.g., Growth: vs. ; Self-Protection: vs. ; Social Focus: vs. ; Personal Focus: vs. ; see Tables 4, 5).
@@ 4.8 Which improvements are statistically robust under fixed compute
First, threshold tuning is a statistically reliable improvement: in every slice with reported tests, tuned thresholds significantly outperform fixed for Direct models (see “Thr. tuning” in Table 6). McNemar analyses show these gains concentrate in subsets of labels (e.g., Universalism/Beneficence in Self-Transcendence; Security/Conformity/Tradition in Conservation; several sparse HO labels), rather than uniformly.
@@ 5 Discussion
(1) HO abstractions are learnable, but they are not uniformly reliable
@@
(2) Calibration and small ensembles are safer bets than hard hierarchies
@@
(3) Presence gating and HO gating improve conditional performance but not end-task performance
@@
(4) Hard gates do not yield reliable end-to-end gains
@@
Under the same budget, prompted and QLoRA-adapted 10B LLMs underperform supervised encoders in absolute Macro-, although few-shot prompting helps. Their main practical benefit is as complementary signals in cross-family ensembles for some slices (Section S7). For instance, Gemma-2-9B-it remains below transformer champions (e.g., Growth vs ), but can still improve cross-family ensembles in selected slices.

### v2

@@ 3.1 Problem formulation and label spaces
We study sentence-level human value detection under Schwartz’s refined theory (Schwartz2012overview). Each sentence may express none, one, or multiple values. Let be a sentence and the binary vector over the 19 basic values.111The benchmark provides attained and constrained annotations per value; we collapse them into a single expressed value signal (Section 3.2).
HO categories.
To test whether coarser abstractions help, we deterministically derive eight HO binary labels from the 19 values following Schwartz2012overview: Openness to Change, Conservation, Personal Focus, Social Focus, Self-Enhancement, Self-Transcendence, Growth, and Self-Protection. Let be the set of HO categories and the basic values grouped under . We define:
@@
Hierarchy representation and consistency.
Let be the binary incidence matrix with
Then Eq. (1) can be written compactly as
where the inequality is interpreted element-wise. A prediction pair is hierarchy-consistent if, for every value and HO category , and imply . This formalizes the structural constraint that a basic value cannot be predicted present while all of its parent HO categories are predicted absent.
Presence label.
@@
Bipolar evaluation slices.
In addition to performance over all eight HO labels, we analyze the four canonical bipolar pairs to expose asymmetries in learnability and error propagation: (i) Openness to Change vs. Conservation, (ii) Self-Enhancement vs. Self-Transcendence, (iii) Personal Focus vs. Social Focus, and (iv) Growth vs. Self-Protection (Schwartz2012overview). For each pair, we report Macro- averaged over the two poles.
@@ 3.2 Dataset and preprocessing
We use the official train/validation/test split (ValueEval24Zenodo; Touche2024) at sentence level. To keep modeling choices unified and compute controlled, we use the benchmark’s English version (machine-translated sentences provided by the benchmark). This yields sentences: train, validation, and test. The split is at the text level; all sentences inherit their source text’s split.
@@ 3.3.1 Direct multi-label prediction (supervised encoder)
The Direct approach follows Yeste2026. Given the pooled sentence representation , we apply a linear layer to produce logits and probabilities . Training minimizes standard multi-label binary cross-entropy: we minimize standard multi-label binary cross-entropy over labels. We do not use class weights, matching Yeste2026.
@@ 3.3.2 CategoryValues hierarchy (HO gating)
Why hard gating can help or hurt.
Hard gating enforces hierarchy consistency by construction, but it also introduces a recall bottleneck. Let denote the event that the ungated value model predicts value correctly on a sentence where , and let denote the event that the parent HO gate is open on that sentence. Under Eq. (5), a gated true positive for value requires both events to occur, so
More conservatively,
Thus, hard gating can improve precision by eliminating structurally inconsistent positives, but it cannot create recall that was absent upstream and may reduce it whenever the parent HO detector misses a true case.
@@ 3.3.3 PresenceCategoryValues cascade
We test a three-stage cascade where Presence acts as a first gate. The Presence formulation follows Yeste2026; here we include the PresenceHOValues cascade as one of the hierarchy-injection strategies under study:
@@
This cascade can improve precision by suppressing spurious positives on non-value sentences, but it can compound errors across stages. In particular, a correct final value prediction requires the conjunction of three events: a correct Presence decision, a correct HO gate decision, and a correct value decision. Therefore, each additional hard gate narrows the feasible prediction space while introducing another potential source of false negatives. We quantify overall and per-pair effects in Section 3.8.
@@ 3.3.4 Instruction-tuned LLM baselines (definition prompting)
We benchmark instruction-tuned open LLMs that fit on a single 8 GB GPU (Llama 3.1 8B, Ministral 8B 2410, Qwen 2.5 7B, Gemma 2 9B). We use the definition-style prompt (best in Yeste2026), presenting one-line definitions for the 19 values from Schwartz2012overview. The model returns only a JSON array of applicable value names.
We evaluate zero-shot and few-shot prompting with in-context examples. Few-shot templates prepend exemplars in the same schema and include at least one null exemplar (empty array) when . We use greedy decoding with max_new_tokens=200. The definition-style template was inherited from the controlled prompt comparison in Yeste2026, where it was the strongest prompt family among the variants tested, so we do not repeat a full prompt-engineering sweep here. The reported LLM results should therefore be read as a comparison under a restricted prompt budget chosen for compute parity and reproducibility, not as an upper bound on achievable LLM performance.
Post-processing and label mapping.
@@ 3.3.5 QLoRA fine-tuning (parameter-efficient LLM adaptation)
We additionally evaluate supervised QLoRA (Dettmers2024), following Yeste2026. Based on validation screening, we fine-tune Gemma 2 9B as the backbone. We train only low-rank adapters (base model frozen) with learning rate and save only adapter weights.
@@ 3.4 Compute-frugal auxiliary signals
Lexicon-derived features. We build sentence vectors from psycholinguistic, affective, moral, and value resources: LIWC-22 (Boyd2022), eMFD (Hopp2021), the Schwartz value lexicon in ValuesML (Kiesel2023), and affective lexica including NRC VAD (Warriner2013; Mohammad2018), NRC EmoLex (MohammadTurney2013), NRC Emotion Intensity (MohammadKiritchenko2018), and WorryWords (Mohammad2024). We aggregate token signals into sentence statistics (counts, relative frequencies, averaged intensities) and standardize using training-set statistics.
Topic features. We attach topic-mixture vectors from unsupervised topic models trained on the training split only: LDA (Blei200), NMF (LeeSeung1999), and BERTopic (Grootendorst2022). At inference, validation/test sentences are mapped to topic vectors using the fixed models.
@@ 3.5 Training protocol and compute parity
Experimental setup at a glance.
All experiments use the official English ValueEval’24/ValuesML train/validation/test split at sentence level. We compare three model families under the same compute budget: supervised DeBERTa-base encoders, prompted instruction-tuned LLMs, and QLoRA-adapted LLMs. For supervised runs, we keep the backbone, maximum sequence length, and optimization protocol fixed across variants, and vary only the decision structure (Direct, CategoryValues, PresenceCategoryValues) and optional low-cost features. Model selection is performed on the validation split only, including threshold tuning and ensemble construction, and all final results are reported once on the held-out test split using the frozen validation-selected configuration.
Unless otherwise stated, we follow the supervised protocol of Yeste2026. All transformer models fine-tune microsoft/deberta-base (He2021) with a linear multi-label head. Inputs are tokenized and truncated/padded to length 512. We optimize with AdamW (LoshchilovHutter2018) and a linear schedule with warmup, using batch size 4, gradient accumulation 4 (effective batch size 16), learning rate , weight decay 0.15, and up to 10 epochs with early stopping on validation Macro- (patience 4). Dropout is 0.1.
@@
To keep the study compute-frugal while covering many strategies (direct prediction, hard gates/cascades, auxiliary signals, prompting, QLoRA, ensembles), we fix a single random seed (as in Yeste2026) for supervised runs. This prioritizes breadth under a fixed budget and keeps differences tied to modeling choices rather than to random initialization. Instead of multiple seeds, we emphasize paired evaluation: nonparametric sentence-level bootstrap uncertainty for Macro- and per-label paired tests (McNemar with FDR correction). We interpret differences below roughly 1–2 Macro- points conservatively and focus on effects supported by paired tests.
@@ 3.6 Threshold calibration
For supervised encoders (and QLoRA models that output probabilities), we use (i) a fixed global threshold or (ii) label-wise thresholds from a constrained grid search over . For each label , we solve
This turns threshold selection into a constrained decision rule: under severe imbalance, we preserve a minimum precision floor while maximizing recall, which is often the limiting factor for Macro- on sparse labels. For hierarchical models, thresholds are tuned in stage-aware order (Presence first, then HO, then values). Prompted LLMs output discrete label sets and do not require threshold calibration.
@@ 3.8 Evaluation metrics and statistical testing
Primary metric.
We report macro-averaged over the label set. We compute per label across all sentences, then average across labels. Macro- is appropriate here because the label distribution is strongly imbalanced and the study aims to evaluate gains on rare as well as frequent values; unlike micro-averaging, it prevents common labels from dominating the summary metric. For bipolar slices (Section 3.1), we compute Macro- for each pole and average the two poles.
End-task evaluation.
@@
Uncertainty and significance.
To assess robustness, we use nonparametric bootstrap resampling over sentences (Efron1979). We draw samples with replacement, recompute Macro-, and estimate (i) a one-sided 95% lower bound for and (ii) a one-sided empirical -value.
Per-label paired tests.
For individual labels, we use McNemar’s test on paired predictions to detect asymmetric error changes (McNemar1947). We correct for multiple comparisons across labels using Benjamini–Hochberg FDR control (BenjaminiHochberg1995) and report corrected significance where relevant.
@@ 3.9 Reproducibility
We log all configurations (preprocessing, model variants, thresholds, ensemble membership) and preserve prediction files for all runs. We also release the value-to-HO mapping and scripts for preprocessing and evaluation to enable exact replication on the benchmark splits (Touche2024). For LLM experiments, we release the prompts/templates and any adapter weights (where redistribution is permitted), along with deterministic decoding settings and post-processing scripts.
@@ 4 Results
Throughout this section, we use three criteria to assess whether a modeling choice is genuinely useful: (i) absolute end-task performance on the full held-out test distribution, (ii) robustness across HO slices rather than isolated wins, and (iii) paired statistical support from bootstrap and McNemar analyses. This distinction is important for hierarchical methods: a component can look strong on a restricted conditional subproblem while still failing to improve the final end-to-end task.
@@ 4.1 Higher-order categories are learnable, but not equally so
HO categories are learnable with compact supervised encoders, but difficulty varies by pair. The easiest pair is Growth vs. Self-Protection (test Macro-, Table S3). Self-Transcendence vs. Self-Enhancement is moderate (best Macro-, Table S9), while Openness vs. Conservation is hardest (best Macro-, Table S7). These patterns track label prevalence: Openness is rare (about 8% vs. about 20% for Conservation, Table A3), which likely drives both lower Macro- and strong pole asymmetry (Openness Conservation in Table 1).
Table 1 summarizes these differences: Growth vs. Self-Protection is most learnable (Macro-), Self-Transcendence vs. Self-Enhancement is moderate (best Macro-), and Openness vs. Conservation remains hardest (best Macro-), with persistent asymmetry (Conservation Openness). The pattern suggests that “constraint/tradition” cues are captured more reliably than “novelty/autonomy” cues at sentence level. We report single-run results and interpret differences under ~1–2 Macro- points cautiously.
@@ 4.4 Presence gating boosts in-gate validation scores but does not robustly improve end-task performance
Table 3 adds a compact sensitivity check for the same hard-gating idea. Keeping the HO model fixed (Baseline) and varying only the Presence gate threshold from to within the same gate family (LIWC22+Ling. Feat.) does not overturn the conclusion: the gated system remains equal to or below the tuned Direct champion in all four pairs, and class-wise HO threshold tuning does not provide a systematic rescue.
| HO pair | Direct best | ||
| fixed / tuned | fixed / tuned | ||
| Growth vs. Self-Protection | 0.580 | 0.570 / 0.580 | 0.570 / 0.560 |
| Social Focus vs. Personal Focus | 0.570 | 0.560 / 0.560 | 0.550 / 0.550 |
| Openness to Change vs. Conservation | 0.420 | 0.412 / 0.384 | 0.416 / 0.414 |
| Self-Transcendence vs. Self-Enhancement | 0.510 | 0.490 / 0.490 | 0.490 / 0.480 |
The most plausible explanation is error compounding: Presence false negatives suppress downstream positives (recall loss), while Presence false positives admit sentences that still look negative to the HO classifier (precision loss). The net effect is that hard gating trades recall for precision in a way that is hard to tune globally. The compact sweep in Table 3 and the full appendix sweeps (Tables S11 and S13–S17) show the same pattern: changing the gate threshold or switching from fixed to tuned HO thresholds produces only minor fluctuations and does not reverse the direct-vs.-gated ranking. This motivates softer alternatives rather than a binary filter.
@@ 4.5 HOvalues hard gating does not translate into reliable gains on fine-grained detection
Under the Growth gate, the best test Macro- for Growth values is about ; under Self-Protection it is about . Section S7 compares the best HO-gated systems against tuned Direct baselines within each HO slice using paired bootstrap tests. The pattern is consistent: hard HO gating does not yield reliable improvements and is sometimes significantly worse, consistent with recall suppression when parent categories are missed. This brittleness is not tied to a single threshold choice: Table 4 already reports the best-performing gated configurations over a range of gate thresholds (from 0.10 to 0.50), yet even these threshold-selected variants do not produce consistent downstream gains. In a sentence-level setting—where signals are short, implicit, and often multi-valued—this loss is hard to recover.
@@ 4.6 Where small instruction-tuned LLMs fit under the same budget
Section S4 benchmarks instruction-tuned 10B LLMs (prompted and QLoRA-adapted) on the same HO-restricted slices. Table 5 reports the best test Macro- for Gemma-2-9B-it under (i) prompting only, (ii) prompting plus a lightweight SBERT gate, and (iii) QLoRA adaptation, and indicates which upgrades are supported by bootstrap tests (Section S7). Overall, even the best LLM variants remain well below the supervised DeBERTa-based champions on test (e.g., Growth: vs. ; Self-Protection: vs. ; Social Focus: vs. ; Personal Focus: vs. ; see Tables 5, 6).
@@ 4.8 Which improvements are statistically robust under fixed compute
First, threshold tuning is a statistically reliable improvement: in every slice with reported tests, tuned thresholds significantly outperform fixed for Direct models (see “Thr. tuning” in Table 7). McNemar analyses show these gains concentrate in subsets of labels (e.g., Universalism/Benevolence in Self-Transcendence; Security/Conformity/Tradition in Conservation; several sparse HO labels), rather than uniformly.
@@
Taken together, these results give a coherent interpretation of the study design and findings. H1 is supported in qualified form: HO categories are learnable, but their usefulness depends strongly on prevalence and lexical concentration. H2 is supported more clearly: hard hierarchical routing improves structural consistency and can inflate conditional scores, yet it does not deliver robust end-task gains because recall losses accumulate across stages. H3 receives the strongest empirical support: under fixed compute, threshold tuning and small ensembles are the only interventions that repeatedly yield statistically supported improvements. The significance of the paper therefore lies not in proposing a new hierarchy method, but in establishing a more precise benchmark-level conclusion: HO structure is most useful as a descriptive or auxiliary inductive bias, whereas the practically reliable improvements come from calibration and lightweight ensembling.
@@ 5 Discussion
The results do more than rank systems: they clarify why some strategies help and others fail under sentence-level sparsity, overlap, and fixed compute. In particular, the experiments separate the representational value of HO abstractions from the decision-theoretic cost of enforcing them as hard gates. This distinction is central to the paper’s novelty framing, because the main contribution is not a single architecture, but a careful empirical analysis of how HO structure should be used in practice under a bounded compute budget.
@@
(1) HO abstractions are learnable, but they are not uniformly reliable.
@@
(2) Calibration and small ensembles are safer bets than hard hierarchies.
@@
(3) Presence gating and HO gating improve conditional performance but not end-task performance.
@@
(4) Hard gates do not yield reliable end-to-end gains.
@@
Under the same budget, prompted and QLoRA-adapted 10B LLMs underperform supervised encoders in absolute Macro-, although few-shot prompting helps. Their main practical benefit is as complementary signals in cross-family ensembles for some slices (Section S7). For instance, Gemma-2-9B-it remains below transformer champions (e.g., Growth vs. ), but can still improve cross-family ensembles in selected slices.